# Rui Mao1, Chao Huang1, Jing Bi1, Zeliang Zhang1, Pooyan Fazli2, Chenliang Xu1†

### v4

@@ 4 Diagnostic Analysis of Factors Affecting MLLMs’ Performance
Table 4: Overall accuracy of LongVA [62] and LongVILA [29] on different #frm ranging from 44 to 128128.
@@
LongVA [62] 336 7B 40.74 43.73 43.32 41.62 39.98 39.39
LongVILA [29] Dynamic 8B 33.00 35.87 35.11 36.46 36.17 36.11
@@
Video-LLaMA2 [9] 336 32 72B 54.15+28.27 71.67+23.89 65.68+36.84 48.52+2.62 59.5+9.09 58.62+24.27
@@
Chat-UniVi [20] 0.65M 25.56 34.44 23.37 25.9 36.36 26.73
VideoChat2 [26] 8 224 7B 2M 39.46+13.9 57.78+23.34 44.84+21.47 58.03+32.13 50.41+14.05 47.01+20.28
Chat-UniVi-v1.5 [20] 1.27M 37.38 47.78 26.53 37.38 46.28 36.11
LongVA [62] 8 336 7B 1.32M 37.54+0.16 51.67+3.89 41.89+15.36 49.51+12.13 56.2+9.92 43.73+7.62
VILA-1.5 [29] 8B 1.21M 35.94 52.78 34.74 42.62 56.2 40.04
Video-LLaMA2.1-AV [9] 7B 3.35M 35.46-0.48 52.78+0 37.05+3.31 39.34-3.28 58.68+2.48 40.09+0.05
Video-LLaMA2.1 [9] 8 384 7B 3.35M 38.34+2.88 57.78+5.0 54.95+17.9 46.89+7.75 48.76-9.92 47.3+7.22
Chat-UniVi [20] 0.65M 24.92 32.22 23.37 24.92 38.84 26.26
VideoChat2 [26] 16 224 7B 2M 39.3+14.38 60.0+27.78 44.84+21.47 31.8+6.88 26.45-12.39 40.8+14.54
Chat-UniVi-v1.5 [20] 1.27M 34.5 48.89 25.89 40.98 42.98 35.4
LongVA [62] 16 336 7B 1.32M 37.86+3.36 51.11+2.22 41.89+16.0 47.87+6.89 53.72+10.74 43.32+7.92
VILA-1.5 [29] 8B 1.21M 35.78 51.11 36.42 39.34 60.33 39.98
Video-LLaMA2.1-AV [9] 7B 3.35M 35.78+0 56.67+5.56 36.42+0 40.0+0.66 59.5-0.83 40.62+0.64
Video-LLaMA2.1 [9] 16 384 7B 3.35M 38.34+2.56 57.78+1.11 54.95+18.53 46.89+6.89 48.76-10.74 47.3+6.68
Kangaroo [31] 2.94M 31.79 51.67 29.05 53.44 33.06 37.1
MiniCPM-V [56] 64 448 8B 8.32M 38.18+6.39 60.0+8.33 40.84+11.79 38.36-15.08 55.37+22.31 42.5+5.4
@@
Qualitative Analysis. We perform an error analysis to gain deeper insights into the models’ shortcomings in fine-grained video composition understanding. In this analysis, the models are required to answer questions and provide explanations in a dialogue format. Figure 5 shows the examples where top models fail to predict correct answers. For example, while humans easily use visual context to distinguish camera movements and angles like “pan left” and “zoom in” or angles like “eye level” and “low angle,” models like LLaVA-OneVision-72B [24] and GPT-4 [1] often struggle due to scene transitions and subtle perspective changes.
@@ 5 Related Work
MLLMs for Video Understanding. Equipping LLMs with adapted video encoders has led to the creation of several multimodal models tailored for video understanding [47]. For instance, GPT-4-turbo, GPT-4o, and GPT-4o-mini [1] are GPT-based models with integrated video comprehension capabilities. InternVL2 [8], with parameter counts ranging from 1B to 76B, is based on the InternLM framework [4] and supports video processing at multiple scales. Additionally, models derived from the LLaMA backbone—such as LLaVA-OneVision [24], VILA [29], VideoLLaMA [9], and LongLLaVA [53]—have been adapted for video input. Gemini has also been extended to include video processing capabilities [44]. Other models, including Qwen [50], MiniCPM [56], Kangaroo [31], and Chat-UniVi [20], exhibit strong video understanding abilities. In our work, we thoroughly evaluate these models’ capabilities in video composition understanding and provide detailed analysis.
Evaluation Benchmarks for MLLMs. Numerous benchmarks for MLLMs have recently emerged to evaluate diverse model capabilities, with image captioning, Visual Question Answering (VQA), and visual reasoning among the most frequently assessed tasks. Image captioning [30, 43, 39, 10] measures an MLLM’s ability to generate text descriptions of visual content. VQA [2, 38, 40] assesses the model’s proficiency in answering questions based on visual inputs by integrating visual perception with language understanding and external knowledge. Visual reasoning [21, 46, 18] evaluates a model’s spatial awareness and logical reasoning in processing visual information. Moreover, the comprehensive abilities of MLLMs are gauged using advanced benchmarks [23, 32, 60, 12, 57, 37, 14, 36, 58]. For video MLLMs, similar efforts leverage existing benchmarks [52, 27] to evaluate video understanding [25, 42]. However, a notable gap remains in evaluating models on the video-composition understanding. This composition understanding is crucial for accurately processing and correlating multiple elements within a visual scene [61, 15]. While existing benchmarks assess compositionally in images [49, 17], few comprehensively address the specific challenges of compositional understanding in video, where many MLLMs still show limitations.

### v5

@@ 4 Diagnostic Analysis of Factors Affecting MLLMs’ Performance
Table 4: Overall accuracy of LongVA [zhang2024longva] and LongVILA [lin2024vila] on different #frm ranging from 44 to 128128.
@@
LongVA [zhang2024longva] 336 7B 40.74 43.73 43.32 41.62 39.98 39.39
LongVILA [lin2024vila] Dynamic 8B 33.00 35.87 35.11 36.46 36.17 36.11
@@
Video-LLaMA2 [cheng2024videollama2] 336 32 72B 54.15+28.27 71.67+23.89 65.68+36.84 48.52+2.62 59.5+9.09 58.62+24.27
@@
Chat-UniVi [jin2024chat-univi] 0.65M 25.56 34.44 23.37 25.9 36.36 26.73
VideoChat2 [li2024mvbench] 8 224 7B 2M 39.46+13.9 57.78+23.34 44.84+21.47 58.03+32.13 50.41+14.05 47.01+20.28
Chat-UniVi-v1.5 [jin2024chat-univi] 1.27M 37.38 47.78 26.53 37.38 46.28 36.11
LongVA [zhang2024longva] 8 336 7B 1.32M 37.54+0.16 51.67+3.89 41.89+15.36 49.51+12.13 56.2+9.92 43.73+7.62
VILA-1.5 [lin2024vila] 8B 1.21M 35.94 52.78 34.74 42.62 56.2 40.04
Video-LLaMA2.1-AV [cheng2024videollama2] 7B 3.35M 35.46-0.48 52.78+0 37.05+3.31 39.34-3.28 58.68+2.48 40.09+0.05
Video-LLaMA2.1 [cheng2024videollama2] 8 384 7B 3.35M 38.34+2.88 57.78+5.0 54.95+17.9 46.89+7.75 48.76-9.92 47.3+7.22
Chat-UniVi [jin2024chat-univi] 0.65M 24.92 32.22 23.37 24.92 38.84 26.26
VideoChat2 [li2024mvbench] 16 224 7B 2M 39.3+14.38 60.0+27.78 44.84+21.47 31.8+6.88 26.45-12.39 40.8+14.54
Chat-UniVi-v1.5 [jin2024chat-univi] 1.27M 34.5 48.89 25.89 40.98 42.98 35.4
LongVA [zhang2024longva] 16 336 7B 1.32M 37.86+3.36 51.11+2.22 41.89+16.0 47.87+6.89 53.72+10.74 43.32+7.92
VILA-1.5 [lin2024vila] 8B 1.21M 35.78 51.11 36.42 39.34 60.33 39.98
Video-LLaMA2.1-AV [cheng2024videollama2] 7B 3.35M 35.78+0 56.67+5.56 36.42+0 40.0+0.66 59.5-0.83 40.62+0.64
Video-LLaMA2.1 [cheng2024videollama2] 16 384 7B 3.35M 38.34+2.56 57.78+1.11 54.95+18.53 46.89+6.89 48.76-10.74 47.3+6.68
Kangaroo [liu2024kangaroo] 2.94M 31.79 51.67 29.05 53.44 33.06 37.1
MiniCPM-V [yao2024minicpm] 64 448 8B 8.32M 38.18+6.39 60.0+8.33 40.84+11.79 38.36-15.08 55.37+22.31 42.5+5.4
@@
Qualitative Analysis. We perform an error analysis to gain deeper insights into the models’ shortcomings in fine-grained video composition understanding. In this analysis, the models are required to answer questions and provide explanations in a dialogue format. Figure 5 shows the examples where top models fail to predict correct answers. For example, while humans easily use visual context to distinguish camera movements and angles like “pan left” and “zoom in” or angles like “eye level” and “low angle,” models like LLaVA-OneVision-72B [li2024llavaonevisioneasyvisualtask] and GPT-4 [achiam2023gpt] often struggle due to scene transitions and subtle perspective changes.
@@ 5 Related Work
MLLMs for Video Understanding. Equipping LLMs with adapted video encoders has led to the creation of several multimodal models tailored for video understanding [tang2023video]. For instance, GPT-4-turbo, GPT-4o, and GPT-4o-mini [achiam2023gpt] are GPT-based models with integrated video comprehension capabilities. InternVL2 [chen2024internvl], with parameter counts ranging from 1B to 76B, is based on the InternLM framework [cai2024internlm2] and supports video processing at multiple scales. Additionally, models derived from the LLaMA backbone—such as LLaVA-OneVision [li2024llavaonevisioneasyvisualtask], VILA [lin2024vila], VideoLLaMA [cheng2024videollama2], and LongLLaVA [wang2024longllava]—have been adapted for video input. Gemini has also been extended to include video processing capabilities [reid2024gemini]. Other models, including Qwen [Qwen2-VL], MiniCPM [yao2024minicpm], Kangaroo [liu2024kangaroo], and Chat-UniVi [jin2024chat-univi], exhibit strong video understanding abilities. In our work, we thoroughly evaluate these models’ capabilities in video composition understanding and provide detailed analysis.
Evaluation Benchmarks for MLLMs. Numerous benchmarks for MLLMs have recently emerged to evaluate diverse model capabilities, with image captioning, Visual Question Answering (VQA), and visual reasoning among the most frequently assessed tasks. Image captioning [lin2014microsoft, OnoeDocci2024, Masry2022ChartQAAB, feng2024more] measures an MLLM’s ability to generate text descriptions of visual content. VQA [antol2015vqa, marino2019ok, Mathew2020DocVQAAD] assesses the model’s proficiency in answering questions based on visual inputs by integrating visual perception with language understanding and external knowledge. Visual reasoning [johnson2017clevr, Suhr2017nlvr, hua2025finematch] evaluates a model’s spatial awareness and logical reasoning in processing visual information. Moreover, the comprehensive abilities of MLLMs are gauged using advanced benchmarks [li2023seed, liu2023mmbench, yue2023mmmu, Fu2023MMEAC, yu2024mm, lu2023mathvista, guan2024hallusionbench, lu2024revisiting, yu2024promptfix]. For video MLLMs, similar efforts leverage existing benchmarks [wang2020vatexlargescalehighqualitymultilingual, li2021valuemultitaskbenchmarkvideoandlanguage] to evaluate video understanding [li2024mvbenchcomprehensivemultimodalvideo, ning2023videobenchcomprehensivebenchmarktoolkit]. However, a notable gap remains in evaluating models on the video-composition understanding. This composition understanding is crucial for accurately processing and correlating multiple elements within a visual scene [yuksekgonul2022and, hsieh2024sugarcrepe]. While existing benchmarks assess compositionally in images [thrush2022winoground, hua2024mmcomposition], few comprehensively address the specific challenges of compositional understanding in video, where many MLLMs still show limitations.